How do i get rid of a coefficent in order to make the equations for ellipses and hyperbolas in standard form? like in the problem: 25(x+4)2 + (y-1)2 = -75
Answer by stanbon(48581)

How do i get rid of a coefficent in order to make the equations for ellipses and hyperbolas in standard form? like in the problem: 25(x+4)2 + (y-1)2 = -75
------------------------
If that 75 is really negative you do not have an ellipse or a hyperpola.
------------------------
If the 75 is positive you get the following:
25(x+4)^2 + (y-1)^2 = 75
Divide by 75 to get:
(x+4)^2/3 + (y-1)^2/75 = 1
=================================
Cheers,
Stan H.